Write a python program that writes to an output file "decoder.txt" a table of letters as follows:

___1_2_3_4_5

1: A B C D E

2: F G H I K

3: L M N O P

4: Q R S T U

5: V W X Y Z

Note that the letter J has been removed.

Use string.ascii_uppercase to start
Answer #1

import string

s = string.ascii_uppercase

j = 0
f = open("decoder.txt","w")
f.write("___1_2_3_4_5\n")
for i in range(5):
   f.write(str(i+1) + ": ")
   # print(str(i+1) + ": ",end='')
   # _ for not using any loop variable
   for _ in range(5):
       if s[j] is 'J':
           j = j+1
       f.write(s[j] + " ")
       # print(s[j],end=" ")
       j = j + 1

f.write("\n")